solve by factoring x(4x-11)=3

Answer by unlockmath(1688)

Hello,
Let's expand this out first.
x(4x-11)=3 rewritten as:
4x^2-11x=3 Subtract 3 from both sides to be:
4x^2-11-3=0 Now we can factor this to give us:
(4x+1)(x-3)=0 Sovingxwillbe:
4x+1=0 Subtract 1 and divide 4 into both sides to result in:
x=-1/4
x=3
This can be by plugging the answers intothe original equation.
RJ
Check out a new book I wrote at:
www.math-unlock.com